# Gravitational Wave Memory from Accelerating Relativistic Jets in Multiple Thick Shell Scenarios

### v1

@@ III.2.1 Single thick shell with common Lorentz factor
We consider in this case the GW memory from a single thick shell expressed by Eq. (14) under situations $N=1$ and $\gamma_{1}=100$. The waveforms of three different thicknesses, where $W_{1}\simeq\delta t_{1}/c$ of a single thick shell, are plotted in Fig. 5 along with the waveform of a single thin shell for comparison. The amplitude does not change at low-frequency because the waveform becomes constant due to $e^{2\pi if\delta t_{j}}\sim 1+i2\pi f\delta t_{j}$ in Eq. (13), which is similar to the discussion in Sec. III.1.2. At high frequencies, attenuation begins under the condition $f\delta t_{j}\geq 1$, such as when $f\sim$ kHz and $\delta t_{j}^{-1}\sim$ kHz. Compared to a single thin shell, high-frequency components around $f\geq$ kHz attenuate more than in the case of a thin shell. This attenuation is due to the effect of thickness in the term $e^{-i2\pi f\delta t_{j}}$ in Eq. (13), and we confirm that the frequency range where the attenuation begins shifts to lower frequencies as the thickness increases.
@@ III.2.2 Spreading of a single thick shell
According to the internal shock model of the prompt GRB emission, the shells spread in the radial direction because the front and rear ends of the shell have different Lorentz factors [1, 26]. As the shells expand, their radial thickness increases. The shell thickness should be sufficiently large ($\sim 10^{10}$ cm) when the collisions of shells occur at $r\sim 10^{13-14}$ cm so that the lifetime of the internal shocks explains the rising time of the observed gamma-ray pulses. Let $\gamma_{\mathrm{front}}$ and $\gamma_{\mathrm{rear}}$ be the Lorentz factors of the front and rear shells, respectively. Suppose that $\gamma_{\mathrm{front}}>\gamma_{\mathrm{rear}}$ and the Lorentz factors within the single thick shell decrease monotonically from front to rear, where $\gamma_{\mathrm{front}}$ and $\gamma_{\mathrm{rear}}$ denote the maximum and minimum values in the shell, respectively. Figure 6 shows the waveform from a single thick shell under this condition. There is no significant difference in the waveform compared to the case that the front and rear are the same Lorentz factor, except for the sharply attenuated peak. This fact is understood as follows. The thickness of a shell just after the end of acceleration is given by:
@@
where $\Delta\gamma=\gamma_{\mathrm{front}}-\gamma_{\mathrm{rear}}$. In the r.h.s. of this equation, the second term is approximately $\sim r_{0}/\gamma_{\mathrm{mid}}\sim 10^{5}~{}\mathrm{cm}$, which is much smaller than the first one, $\sim 10^{7}~{}\mathrm{cm}$. Therefore, in the remainder of this paper, we neglect the effect of the shell spreading.
Figure 6: The waveform when Lorentz factor at the front of a single thick shell is larger than that at the rear. The Lorentz factors of the front and rear shells follow the condition $\gamma_{\mathrm{front}}>\gamma_{\mathrm{rear}}$, which represents a thick shell spreading over time. The dashed black line shows the waveform having the same Lorentz factor for both front and rear shells. The solid blue line in Fig. 5 is used as a reference. We adopt $\mathcal{E}=10^{52}~{}\mathrm{erg}$, $W=10^{7}~{}\mathrm{cm}$, $R=10~{}\mathrm{Mpc}$, $\gamma_{1}$ = 100, and $\Delta\theta=\theta_{v}=0.1~{}\mathrm{rad}$.
@@
The waveform formula from multiple thin shells is expressed in Eq. (11). We plot this waveforms for shell numbers $N=1,2,10,\mathrm{and}~{}100$ in Fig. 7. Each waveform is configured where the $N$ shells are launched with equal time intervals $\delta\tau_{j}=0.1$ sec and with the same Lorentz factor $\gamma_{j}=100$ for all $j$, and $m_{j}$ in Eq. (9) as constant. For comparison, The total energy of each case is set to constant as $\mathcal{E}=10^{52}~{}\mathrm{erg}$. The amplitude attenuation begins at low-frequency under the condition $fT_{\mathrm{CE}}\geq 1$, such as $f\sim 100$ Hz and $T_{\mathrm{CE}}^{-1}\sim 100$ Hz in the case of $N=100$. As a result of the time shift of the Fourier transform in Eq. (11), this impact appears at lower frequencies as the number of shells increases. Therefore, this attenuation could help in understanding how many shells are formed in the central engine of the jet.
@@
Figure 7: Characteristic strain of GW from multiple thin shells. $N$ represents the number of thin shells. $N$ shells are emitted with equal time intervals, $\delta\tau_{j}=0.1$ for all $j$. The total time is $T_{\mathrm{CE}}=0.1$ sec in the case of $N=2$, $T_{\mathrm{CE}}\sim 1$ sec in the case $N=10$, and $T_{\mathrm{CE}}\sim 10$ sec in the case of $N=100$. The dashed black line is plotted as a reference, which is same as the solid purple line in Fig. 3. All plots have $\mathcal{E}=10^{52}~{}\mathrm{erg}$, $R=10~{}\mathrm{Mpc}$, $\gamma_{j}$ = 100 for all $j$, and $\Delta\theta=\theta_{v}=0.1~{}\mathrm{rad}$.
@@ III.4.1 Constant $\gamma_{j}$ and $\delta\tau_{j}$
We compute the waveform with uniform time interval $\delta\tau_{j}=0.1$ rad and Lorentz factor $\gamma_{j}=100$ for all $j$. Figure 8 shows the waveforms from $N$ thick shells along with $N=2$ thin shells for comparison. The amplitude attenuation appears at high-frequency, similarly to the discussion of Fig. 5 in Sec. III.2. Such attenuation also appears at lower frequencies as the number of shells increases, which is the same feature as shown in Fig. 7. Waveforms with different thicknesses are shown in Fig. 9. In the same way as the discussion of Fig. 5, the frequency range where the attenuation begins shifts to lower frequencies as the thickness increases. We confirm that the effect of shell thickness and the number of shells appears independently in the waveform.
Figure 8: (a) Top: Multiple thick shells model waveform, with $N=2$ being the number of shells, and the corresponding thin shell model waveform in dashed gray as a comparison. (b) Bottom: Multiple thick shells, with $N=10$, and the corresponding thin shell model waveform in dashed gray as a comparison. All plots have $\mathcal{E}=10^{52}~{}\mathrm{erg}$, $W=10^{7}\mathrm{cm}$, $R=10~{}\mathrm{Mpc}$, $\gamma_{j}$ = 100 for all $j$, $\delta\tau_{j}=0.1$ and $\Delta\theta=\theta_{v}=0.1~{}\mathrm{rad}$.
Figure 9: (a) Top: Multiple thick shells model waveform, with a thickness $W=10^{8}~{}\mathrm{cm}$, and the corresponding thin shell model waveform in dashed gray with $W=10^{7}~{}\mathrm{cm}$ as a comparison. (b) Bottom: Multiple thick shells with $W=10^{9}~{}\mathrm{cm}$ and the corresponding thin shell model waveform in dashed gray as a comparison. All plots have $\mathcal{E}=10^{52}~{}\mathrm{erg}$, $N=2$, $R=10~{}\mathrm{Mpc}$, $\gamma_{j}$ = 100 for all $j$, $\delta\tau_{j}=0.1$ and $\Delta\theta=\theta_{v}=0.1~{}\mathrm{rad}$.
@@ III.4.2 Single parameter randomization: either $\gamma_{j}$ or $\delta\tau_{j}$
At the first stage and for simplicity purposes, we vary $\gamma_{j}$ or random $\delta\tau_{j}$ while keeping the other fixed to observe their individual effects on the waveform. We consider the waveforms from multiple thick shells with a random $\gamma_{j}\in[100,1000]$ and a constant $\delta\tau_{j}=0.1$, shown in Fig. 12. Figure 12 shows the $\gamma_{j}$ assigned to each shell at each time. The quantile–quantile (Q–Q) plot, which compares the samples to a specific distribution, indicates the sampled $\gamma_{j}$ follows a uniform distribution. The solid gray line represents a constant $\gamma_{j}=500$, with the total energy adjusted to be equivalent to that in the case of random $\gamma_{j}$. Due to differences in $\gamma_{j}$, slight variations in amplitude can be seen in the frequency range $f\in[10^{0},10^{2}]$ Hz. In actual GRBs, constant $\gamma_{j}$ would prevent interactions between shells, leading to an expected waveform similar to the solid red line, while the 10 Hz peak remains visible because $\delta\tau_{j}$ is constant.
The waveform with constant $\gamma_{j}=100$ and random $\delta\tau_{j}$ is shown in Fig. 12. The time intervals $\delta\tau_{j}$ are randomized in about 10 seconds, as shown in Fig. 12. Due to the randomness of the time intervals, the frequency peak predicted by Eq. (18) is no longer present, resulting in a flattened waveform in the frequency range $f\in[10^{1},10^{3}]$ Hz.
@@
Figure 10: (a) Left: Waveform under the random $\gamma_{j}$. The solid gray line represents the waveform with $\gamma_{j}=500$, which is the median value for the random $\gamma_{j}$, shown here as a reference. The corresponding energies are $\mathcal{E}=5.75\times 10^{52}~{}\mathrm{erg}$ (red) and $5\times 10^{52}~{}\mathrm{erg}$ (gray). All plots have $W=10^{7}~{}\mathrm{cm}$, $N=100$, $R=10~{}\mathrm{Mpc}$, $\delta\tau_{j}=0.1$ and $\Delta\theta=\theta_{v}=0.1~{}\mathrm{rad}$.
@@
(b) Right: The random $\gamma_{j}\in[100,1000]$ and random $\delta\tau_{j}$ are applied for each shell, and $T_{\mathrm{CE}}=9.91$ sec. A $p$-value from $\chi^{2}$ test is 0.911, assessing whether $\tau_{j}$ and $\gamma_{j}$ are independent.
@@ III.4.3 Randomization $\gamma_{j}$ and $\delta\tau_{j}$
In this case, the Lorentz factor $\gamma_{j}$ and the time interval $\delta\tau_{j}$ of the emitted shells are both randomized to reflect a more realistic scenario. Figure 12 displays the resulting waveform from multiple thick shells with randomized values of $\gamma_{j}$ and $\delta\tau_{j}$, which values and corresponding Q–Q plots are plotted in Fig. 12. A $p$-value from the $\chi^{2}$ test assesses whether $\tau_{j}$ and $\gamma_{j}$ are independent, which are sampled from a uniform distribution. The result value indicates insufficient evidence to reject the null hypothesis of independence. Therefore, we confirm that there is no correlation between them. The waveform has a flattened shape in the frequency range around 10 Hz to 100 Hz due to the randomness $\delta\tau_{j}$, which is the same feature shown in Fig. 12. However, the randomness $\gamma_{j}$ doesn’t affect the waveform much, and it just contributes to the total energy expressed in Eq. (16).
We demonstrate the waveform implementations of GW memory under various configurations: examining the dependency on the opening half-angle $\Delta\theta$, the viewing angle $\theta_{v}$, and the initial radius $r_{0}$, waveforms from multiple thin and thick shells, and variations in final Lorentz factors and time intervals. Distinct waveform shapes are confirmed across different setups.
IV Detectability in near-real scenarios
In this section, we examine the detectability of more realistic GW memory by the next-generation detectors, ET and DECIGO. The detector overviews are described in the Appendix A. As part of this realistic scenario, both the terminal Lorentz factor $\gamma_{j}$ and the time interval of emitted shells $\delta\tau_{j}$ are supposed to be random.
@@ IV.1 Waveforms with typical $T_{\mathrm{CE}}$
Based on actual GRBs observations [35], e.g., in the Burst And Transient Source Experiment (BATSE) catalog [36], the value of $T_{\mathrm{CE}}$ ranges from a few seconds to hundred seconds. Therefore, we compute waveforms with different $T_{\mathrm{CE}}$ shown in Fig. 13 along with the DECIGO and ET-D sensitivity curves. All waveforms are computed with randomly assigned $\gamma_{j}$ and $\delta\tau_{j}$ parameters to better reflect realistic conditions. The waveforms lie on DECIGO’s sensitivity curve in the low-frequency range, on ET-D’s curve in the high-frequency range, and around 10 Hz, it overlaps with both sensitivity curves. This result suggests the possibility of simultaneously observing the GW memory from GRBs with both ET-D and DECIGO. The longer the $T_{\mathrm{CE}}$, the more the initial attenuation shifts toward the lower frequency band, which band would be possible to observable in the $\leq 1$ Hz range by DECIGO. As discussed in Sec. III, the attenuation amplitude over the frequency range depends on the physical configuration of the GRBs. Therefore, a key step toward further understanding GRBs is to observe their GW memories using multiple detectors that cover a wide range of frequency bands.

### v2

@@ III.2.1 Single thick shell with common Lorentz factor
We consider in this case the GW memory from a single thick shell expressed by Eq. (14) under situations $N=1$ and $\gamma_{1}=100$. The waveforms of three different thicknesses, where $W_{1}\simeq\delta t_{1}/c$ of a single thick shell, are plotted in Fig. 5 along with the waveform of a single thin shell for comparison. The amplitude does not change at low frequency because the waveform becomes constant due to $e^{2\pi if\delta t_{j}}\sim 1+i2\pi f\delta t_{j}$ in Eq. (13), which is similar to the discussion in Sec. III.1.2. At high frequencies, attenuation begins under the condition $f\delta t_{j}\geq 1$, such as when $f\sim$ kHz and $\delta t_{j}^{-1}\sim$ kHz. Compared to a single thin shell, high-frequency components around $f\geq$ kHz attenuate more than in the case of a thin shell. This attenuation is due to the effect of thickness in the term $e^{-i2\pi f\delta t_{j}}$ in Eq. (13), and we confirm that the frequency range where the attenuation begins shifts to lower frequencies as the thickness increases.
@@ III.2.2 Spreading of a single thick shell
According to the internal shock model of the prompt GRB emission, the shells spread in the radial direction because the front and rear ends of the shell have different Lorentz factors [1, 27]. As the shells expand, their radial thickness increases. The shell thickness should be sufficiently large ($\sim 10^{10}$ cm) when the collisions of shells occur at $r\sim 10^{13-14}$ cm so that the lifetime of the internal shocks explains the rising time of the observed gamma-ray pulses. Let $\gamma_{\mathrm{front}}$ and $\gamma_{\mathrm{rear}}$ be the Lorentz factors of the front and rear shells, respectively. Suppose that $\gamma_{\mathrm{front}}>\gamma_{\mathrm{rear}}$ and the Lorentz factors within the single thick shell decrease monotonically from front to rear, where $\gamma_{\mathrm{front}}$ and $\gamma_{\mathrm{rear}}$ denote the maximum and minimum values in the shell, respectively. Figure 6 shows the waveform from a single thick shell under this condition. There is no significant difference in the waveform compared to the case that the front and rear are the same Lorentz factor, except for the sharply attenuated peak. This fact is understood as follows. The thickness of a shell just after the end of acceleration is given by
@@
where $\Delta\gamma=\gamma_{\mathrm{front}}-\gamma_{\mathrm{rear}}$. In the rhs of this equation, the second term is approximately $\sim r_{0}/\gamma_{\mathrm{mid}}\sim 10^{5}~\mathrm{cm}$, which is much smaller than the first one, $\sim 10^{7}~\mathrm{cm}$. Therefore, in the remainder of this paper, we neglect the effect of the shell spreading.
Figure 6: The waveform when the Lorentz factor at the front of a single thick shell is larger than that at the rear. The Lorentz factors of the front and rear shells follow the condition $\gamma_{\mathrm{front}}>\gamma_{\mathrm{rear}}$, which represents a thick shell spreading over time. The dashed black line shows the waveform having the same Lorentz factor for both front and rear shells. The solid blue line in Fig. 5 is used as a reference. We adopt $\mathcal{E}=10^{52}~\mathrm{erg}$, $W=10^{7}~\mathrm{cm}$, $R=10~\mathrm{Mpc}$, $\gamma_{1}$ = 100, and $\Delta\theta=\theta_{v}=0.1~\mathrm{rad}$.
@@
The waveform formula from multiple thin shells is expressed in Eq. (11). We plot this waveform for shell numbers $N=1,2,10,\mathrm{and}~100$ in Fig. 7. Each waveform is configured where the $N$ shells are launched with equal time intervals $\delta\tau_{j}=0.1$ sec and with the same Lorentz factor $\gamma_{j}=100$ for all $j$, and $m_{j}$ in Eq. (9) as constant. For comparison, the total energy of each case is set to constant as $\mathcal{E}=10^{52}~\mathrm{erg}$. The amplitude attenuation begins at low frequency under the condition $fT_{\mathrm{CE}}\geq 1$, such as $f\sim 100$ Hz and $T_{\mathrm{CE}}^{-1}\sim 100$ Hz in the case of $N=100$. As a result of the time shift of the Fourier transform in Eq. (11), this impact appears at lower frequencies as the number of shells increases. Therefore, this attenuation could help in understanding how many shells are formed in the central engine of the jet.
@@
Figure 7: Characteristic strain of GW from multiple thin shells. NN represents the number of thin shells. NN shells are emitted with equal time intervals, $\delta\tau_{j}=0.1$ for all $j$. The total time is $T_{\mathrm{CE}}=0.1$ sec in the case of $N=2$, $T_{\mathrm{CE}}\sim 1$ sec in the case of $N=10$, and $T_{\mathrm{CE}}\sim 10$ sec in the case of $N=100$. The dashed black line is plotted as a reference, which is the same as the solid purple line in Fig. 3. All plots have $\mathcal{E}=10^{52}~\mathrm{erg}$, $R=10~\mathrm{Mpc}$, $\gamma_{j}$ = 100 for all $j$, and $\Delta\theta=\theta_{v}=0.1~\mathrm{rad}$.
@@ III.4.1 Constant $\gamma_{j}$ and $\delta\tau_{j}$
We compute the waveform with uniform time interval $\delta\tau_{j}=0.1$ rad and Lorentz factor $\gamma_{j}=100$ for all $j$. Figure 8 shows the waveforms from $N$ thick shells along with $N=2$ thin shells for comparison. The amplitude attenuation appears at high frequency, similarly to the discussion of Fig. 5 in Sec. III.2. Such attenuation also appears at lower frequencies as the number of shells increases, which is the same feature as shown in Fig. 7. Waveforms with different thicknesses are shown in Fig. 9. In the same way as the discussion of Fig. 5, the frequency range where the attenuation begins shifts to lower frequencies as the thickness increases. We confirm that the effect of shell thickness and the number of shells appear independently in the waveform.
Figure 8: (a) Top: multiple thick shell model waveform, with $N=2$ being the number of shells, and the corresponding thin shell model waveform in dashed gray as a comparison. (b) Bottom: multiple thick shells, with $N=10$, and the corresponding thin shell model waveform in dashed gray as a comparison. All plots have $\mathcal{E}=10^{52}~\mathrm{erg}$, $W=10^{7}\mathrm{cm}$, $R=10~\mathrm{Mpc}$, $\gamma_{j}$ = 100 for all $j$, $\delta\tau_{j}=0.1$ and $\Delta\theta=\theta_{v}=0.1~\mathrm{rad}$.
Figure 9: (a) Top: multiple thick shell model waveform, with a thickness $W=10^{8}~\mathrm{cm}$, and the corresponding thin shell model waveform in dashed gray with $W=10^{7}~\mathrm{cm}$ as a comparison. (b) Bottom: multiple thick shells with $W=10^{9}~\mathrm{cm}$ and the corresponding thin shell model waveform in dashed gray as a comparison. All plots have $\mathcal{E}=10^{52}~\mathrm{erg}$, $N=2$, $R=10~\mathrm{Mpc}$, $\gamma_{j}$ = 100 for all $j$, $\delta\tau_{j}=0.1$ and $\Delta\theta=\theta_{v}=0.1~\mathrm{rad}$.
@@ III.4.2 Single parameter randomization: Either $\gamma_{j}$ or $\delta\tau_{j}$
At the first stage and for simplicity purposes, we vary $\gamma_{j}$ or random $\delta\tau_{j}$ while keeping the other fixed to observe their individual effects on the waveform. We consider the waveforms from multiple thick shells with a random $\gamma_{j}\in[100,1000]$ and a constant $\delta\tau_{j}=0.1$, shown in Fig. 12. Figure 12 shows the $\gamma_{j}$ assigned to each shell at each time. The quantile–quantile (Q–Q) plot, which compares the samples to a specific distribution, indicates the sampled $\gamma_{j}$ follows a uniform distribution. The solid gray line represents a constant $\gamma_{j}=500$, with the total energy adjusted to be equivalent to that in the case of random $\gamma_{j}$. Owing to differences in $\gamma_{j}$, slight variations in amplitude can be seen in the frequency range $f\in[10^{0},10^{2}]$ Hz. In actual GRBs, constant $\gamma_{j}$ would prevent interactions between shells, leading to an expected waveform similar to the solid red line, while the 10 Hz peak remains visible because $\delta\tau_{j}$ is constant.
The waveform with constant $\gamma_{j}=100$ and random $\delta\tau_{j}$ is shown in Fig. 12. The time intervals $\delta\tau_{j}$ are randomized in about 10 seconds, as shown in Fig. 12. Owing to the randomness of the time intervals, the frequency peak predicted by Eq. (18) is no longer present, resulting in a flattened waveform in the frequency range $f\in[10^{1},10^{3}]$ Hz.
@@
Figure 10: (a) Left: waveform under the random $\gamma_{j}$. The solid gray line represents the waveform with $\gamma_{j}=500$, which is the median value for the random $\gamma_{j}$, shown here as a reference. The corresponding energies are $\mathcal{E}=5.75\times 10^{52}~\mathrm{erg}$ (red) and $5\times 10^{52}~\mathrm{erg}$ (gray). All plots have $W=10^{7}~\mathrm{cm}$, $N=100$, $R=10~\mathrm{Mpc}$, $\delta\tau_{j}=0.1$, and $\Delta\theta=\theta_{v}=0.1~\mathrm{rad}$.
@@
(b) Right: the random $\gamma_{j}\in[100,1000]$ and random $\delta\tau_{j}$ are applied for each shell, and $T_{\mathrm{CE}}=9.91$ sec. A $p$ value from $\chi^{2}$ test is 0.911, assessing whether $\tau_{j}$ and $\gamma_{j}$ are independent.
@@ III.4.3 Randomization $\gamma_{j}$ and $\delta\tau_{j}$
In this case, the Lorentz factor $\gamma_{j}$ and the time interval $\delta\tau_{j}$ of the emitted shells are both randomized to reflect a more realistic scenario. Figure 12 displays the resulting waveform from multiple thick shells with randomized values of $\gamma_{j}$ and $\delta\tau_{j}$, whose values and corresponding Q–Q plots are plotted in Fig. 12. A $p$ value from the $\chi^{2}$ test assesses whether $\tau_{j}$ and $\gamma_{j}$ are independent, which are sampled from a uniform distribution. The result value indicates insufficient evidence to reject the null hypothesis of independence. Therefore, we confirm that there is no correlation between them. The waveform has a flattened shape in the frequency range around 10 Hz to 100 Hz due to the randomness $\delta\tau_{j}$, which is the same feature shown in Fig. 12. However, the randomness $\gamma_{j}$ does not affect the waveform much, and it just contributes to the total energy expressed in Eq. (16).
We demonstrate the waveform implementations of GW memory under various configurations: examining the dependency on the opening half angle $\Delta\theta$, the viewing angle $\theta_{v}$, and the initial radius $r_{0}$, waveforms from multiple thin and thick shells, and variations in final Lorentz factors and time intervals. Distinct waveform shapes are confirmed across different setups.
IV Detectability in quasirealistic scenarios
In this section, we examine the detectability of more realistic GW memory by the next-generation detectors, ET and DECIGO. The detector overviews are described in Appendix A. As part of this realistic scenario, both the terminal Lorentz factor $\gamma_{j}$ and the time interval of emitted shells $\delta\tau_{j}$ are supposed to be random.
@@ IV.1 Waveforms with typical $T_{\mathrm{CE}}$
Based on actual GRB observations [37], e.g., in the Burst And Transient Source Experiment (BATSE) catalog [38], the value of $T_{\mathrm{CE}}$ ranges from a few seconds to a few hundred seconds. Therefore, we compute waveforms with different $T_{\mathrm{CE}}$ shown in Fig. 13 along with the DECIGO and ET-D sensitivity curves. All waveforms are computed with randomly assigned $\gamma_{j}$ and $\delta\tau_{j}$ parameters to better reflect realistic conditions. The waveforms lie on DECIGO’s sensitivity curve in the low-frequency range, on ET-D’s curve in the high-frequency range, and around 10 Hz, it overlaps with both sensitivity curves. This result suggests the possibility of simultaneously observing the GW memory from GRBs with both ET-D and DECIGO. The longer the $T_{\mathrm{CE}}$, the more the initial attenuation shifts toward the lower frequency band, which would be observable in the $\leq 1$ Hz range by DECIGO. As discussed in Sec. III, the attenuation amplitude over the frequency range depends on the physical configuration of the GRBs. Therefore, a key step toward further understanding of GRBs is to observe their GW memories using multiple detectors that cover a wide range of frequency bands.